OpenStudy (anonymous):

Find the slope of the line that contains the points named. K(c + d, c - d), L(c - d, c + d) -1 1

ganeshie8 (ganeshie8):

use the slope formula ?

ganeshie8 (ganeshie8):

slope of line passong through points \((x_1, y_1)\) and \((x_2, y_2)\) is : \(\large \frac{y_2 - y_1}{x_2-x_1}\)

ganeshie8 (ganeshie8):

K(c + d, c - d), L(c - d, c + d) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plug them above

OpenStudy (anonymous):

so positive two ????????? right?

ganeshie8 (ganeshie8):

did u get positive two ? it doesnt look correct :o

OpenStudy (anonymous):

I mean psitive 1

OpenStudy (anonymous):

*positive

ganeshie8 (ganeshie8):

K(c + d, c - d), L(c - d, c + d) x1 y1 x2 y2 \(\large \frac{y_2 - y_1}{x_2-x_1}\) \(\large \frac{(c+d) - (c-d)}{(c-d)-(c+d)}\)

ganeshie8 (ganeshie8):

K(c + d, c - d), L(c - d, c + d) x1 y1 x2 y2 \(\large \frac{y_2 - y_1}{x_2-x_1}\) \(\large \frac{(c+d) - (c-d)}{(c-d)-(c+d)}\) \(\large \frac{c+d - c+d}{c-d-c-d}\) \(\large \frac{2d}{-2d}\) \(\large -1\)

OpenStudy (anonymous):

thanks ill give you a medal for being so patient with me lol

ganeshie8 (ganeshie8):

np =))